### v1

# for 3D medical image understanding Technical report

Abstract
Vision–language pre-training, i.e., aligning images with paired text, is a powerful paradigm to create encoders that can be directly used for tasks such as classification and retrieval, and for downstream tasks such as segmentation and report generation. In the 3D medical image domain, these capabilities allow vision–language encoders to support radiologists by retrieving patients with similar abnormalities or predicting likelihoods of abnormality. While the methodology holds promise, data availability limits the capabilities of current 3D VLEs.
In this paper, we alleviate the lack of data by injecting additional inductive biases: introducing a report generation objective and pairing vision–language pre-training with vision-only pre-training. This allows us to leverage both image-only and paired image–text 3D datasets, increasing the total amount of data to which our model is exposed. Through these additional inductive biases, paired with best practices of the 3D medical imaging domain, we develop the comprehensive language-image pre-trained (COLIPRI) encoder family. Our COLIPRI encoders achieve state-of-the-art performance in report generation, classification probing, and zero-shot classification, and remain competitive for semantic segmentation.
AUROC
area under the receiver operating characteristic curve
AUPRC
area under the precision-recall curve
CLIP
contrastive language–image pre-training
CLM
causal language modelling
COLIPRI
comprehensive language–image pre-training
CXR
chest X-ray
DSC
Dice similarity coefficient
EHR
electronic health record
FOV
field of view
LLM
large language model
MAE
masked autoencoder
MIM
masked image modelling
MLP
multilayer perceptron
MSD
Medical Segmentation Decathlon
NLST
National Lung Screening Trial
NSD
normalised surface distance
PCA
principal component analysis
PHI
protected health information
RRG
radiology report generation
SSL
self-supervised learning
SOTA
state-of-the-art
ViT
vision transformer
VLE
vision–language encoder
VLM
vision–language model
VQA
visual question answering
1 Introduction
[…]
CLIP [31] has established itself as one of the strongest paradigms to learn general-purpose image and text representations. Aside from being a potent starting point for adaptation to downstream tasks of interest [11, 23], having language-aligned vision embeddings allows leveraging natural language for open-set classification [31] and open-set segmentation [55].
In 3D medical imaging, this training paradigm is particularly relevant because i) a clinician’s report typically accompanies every image acquired in a clinical setting. Such paired data is therefore abundant within hospitals, even if, due to privacy concerns, reports are rarely publicly shared. Moreover, ii) the contrastive language–image pre-training (CLIP) objective aligns a global representation of the image with an associated report, which enables multimodal retrieval (text-to-image and image-to-text) using learnt latent representations. This semantic search can provide radiologists with a valuable set of reference cases that can help guide treatment decisions or serve as an educational tool. Additionally, iii) the zero-shot classification capabilities can support clinical decision making by providing a fast and cheap first opinion [52], while iv) the zero-shot segmentation provides a way to ground the decision on the scan. This has the potential to allow a clinician to quickly validate or discard the proposal made by the model.
Figure 1: Out COLIPRI encoders yield general-purpose embeddings that can be adapted to a plethora of tasks, reaching state-of-the-art performance in multiple downstream tasks.
Despite these promises, the field of vision–language pre-training with medical images is not as mature as its general-domain counterpart. While CLIP [31], Perception Encoder [5] or SigLIP 2 [43] are well established in the general domain, 3D medical VLEs have only recently started to garner attention [12, 3]. We believe this can be attributed to two key issues: 1) the lack of large, publicly available datasets in the medical domain and 2) domain and modality-specific methodological and engineering hurdles. In this work, we address the above issues by demonstrating how to successfully adapt a vision–language pre-training approach to the 3D medical imaging domain, using image-only and image–text open-access CT datasets.
Data availability
The only currently available large-scale 3D image–report pair datasets are CT-RATE, (25k image–report pairs) [12], INSPECT (19k image-report pairs) [15] [^1] INSPECT only holds Impression sections and not Findings., BIMCV-R (8k image-report pairs) [8] and the dataset of Merlin (25.5k image–report pairs) [3]. While these dataset sizes are substantial within the field of 3D imaging, their combined data scale of about 78k image–report pairs is hugely far from the 400 million image–text pairs that the first CLIP model [31] was trained on, and even further from the scale of the WebLI dataset’s 10 billion images and 12 billion alt-texts [7] used in SigLIP 2 [43]. Aside from report-paired public datasets, the large number of available unpaired images has the potential to substantially increase the amount of overall usable data, with singular datasets like UK Biobank [20] containing more than 100k full-body MRIs, the National Lung Screening Trial (NLST) dataset [39] containing 73k different chest CTs, and the OpenMind dataset [44] containing 114k 3D brain MRI volumes. [^2] Due to focusing on Chest-CT, we only leverage NLST in this study. Although most of these 3D studies were likely acquired with corresponding clinical reports, such reports are not publicly released, resulting in image-only datasets being far more abundant than paired image–report data.
Engineering challenges
The number of voxels in 3D medical images is orders of magnitude larger than the number of pixels in images from the general domain and 2D medical images. For example, a typical chest CT volume may be composed of $\times512$$\times$200 voxels, which makes using entire images in native resolution during training challenging due to the excessive VRAM requirements. A whole-body CT scan would be even larger, with over a thousand axial slices acquired. Subsequently, it is common to either train with crops of images [34, 29] or downsampled images [12, 3]. While the former solution may complicate CLIP-style training as clinical reports refer to the entire volume rather than just the field of view (FOV) of the sub-crop, the latter discards image information, which may be crucial for detecting specific abnormalities.
Key contributions
In this work, we improve the current state of the art in 3D medical vision–language models by leveraging best practices of the 3D medical imaging domain and introducing various inductive biases aimed at making the most of the available data. Our contributions can be summarised as follows:
  1. 1.
    We investigate key design choices of the CLIP training paradigm in the context of 3D medical imaging to extract the maximal value from the limited available vision–language data.
  2. 2.
    We increase the supervision gained from the comprehensive text report by introducing a radiology report generation (RRG) objective akin to CapPa [42].
  3. 3.
    We introduce a vision-only self-supervised objective in conjunction with the CLIP objective, akin to Naeem et al. [26], Maninis et al. [24], allowing us to include unpaired data into the training set and adding a more localised objective for dense downstream tasks.
We evaluate the resulting models holistically through zero-shot classification, classification probes, report generation, and semantic segmentation (Fig. 1), highlighting the strengths and limitations of our current encoders.
[…]
2 Related work
Pre-training in natural imaging
In natural imaging, training from scratch has long been outperformed by leveraging pre-trained encoders as a starting point. Initial works leveraged supervised pre-training like Big Transfer [17], which has been slowly superseded by more advanced self-supervised approaches that operate at both the global and patch level. Among these, iBOT [56] introduced a MIM paradigm that enforces similarity between the features generated by a student network and a teacher network, both at a local and global patch level, demonstrating strong transfer to global and dense imaging tasks. Following this, DINOv2 [27] optimised the iBOT clustering, scaled data, and improved data curation, generating the first general-purpose features that exceeded OpenCLIP [9].
In parallel, CLIP [31] was introduced and gained popularity due to its emergent zero-shot classification capabilities and powerful global understanding performance. Follow-ups i) introduced a sigmoid loss that scales favourably to larger batch sizes (SigLIP) [53]. ii) Showed that captioning-based pre-training (CapPa) [42] can match or exceed contrastive pre-training on various visual question answering (VQA) and classification tasks. iii) Introduced localised captions (LocCa) [47] in the generative objective, showing improved performance on dense tasks. Moreover, iv) Perception Encoder [5] proposes various dataset curation, regularisation, and augmentation changes, among others, to optimise the original CLIP paradigm. Collectively, weakly-supervised vision–language pre-training has been shown to excel at global tasks and image retrieval, while patch-level self-supervised methods like DINOv2 tend to perform better on dense tasks.
Recent works have merged these two paradigms by introducing patch-based self-supervised objectives into a CLIP training framework [26, 24], improving performance on dense downstream tasks, with the recent SigLIP 2 [43] combining CLIP, generative captioning, and self-supervised objectives. Orthogonally, AM-RADIO [32] merges a variety of vision foundation models through a multi-teacher distillation approach, hoping to learn embeddings that excel at both global and local downstream tasks. Recently, Siméoni et al. [35] proposed DINOv3, which achieved state-of-the-art performance on dense and global tasks with self-supervised learning (SSL) alone by scaling DINOv2 and leveraging Gram Anchoring to maintain early dense task performance later in training.
Pre-training in 3D medical imaging
Compared to natural imaging, 3D medical imaging incurs much higher computational costs due to its higher dimensionality and faces tighter data access constraints, limiting the overall amount of available data. These factors have slowed the progress in achieving general-purpose 3D encoders, as most works [57, 37, 49] pre-trained on different small-scale datasets and use various non-state-of-the-art architectures [46].
Nonetheless, recent efforts make large-scale datasets available, [25, 44], and benchmark the currently available SSL strategies of the 3D imaging domain at scale. Across these evaluations, masked autoencoder (MAE[13, 25, 46] proved to be the currently strongest dense pre-training baseline for volumetric segmentation, while contrastive pre-training schemes [38, 50] proved superior for global tasks. However, no pre-training method has shown so far to deliver good performance in dense and global downstream tasks.
Vision–language pre-training in 3D
Analogous to the 3D SSL domain, 3D vision–language model development has been limited by the available datasets. With the recent publication of BIMCV-R [8] and CT-RATE [12], interest in 3D VLEs has increased. The earliest work, CT-CLIP [12], transferred the default CLIP paradigm to 3D; BIUD [6] leveraged existing chest X-rays knowledge to improve CT understanding; Merlin [3] used abdominal CT image-report pairs with electronic health records diagnosis codes as an additional supervisory signal. More recently, fVLM [33] went beyond global image–report alignment, introducing anatomy-wise fine-grained alignment. They used precomputed TotalSegmentator [48] organ masks, decomposed reports into anatomy-specific snippets, and contrasted aligned region-sentence pairs while correcting false negatives, yielding sizeable AUROC gains over CT-CLIP and other benchmarks.
Despite increased interest in the field of 3D self-supervised and vision–language encoders, a research gap remains in methodological advancement between natural imaging and 3D medical imaging. Thus far, no work has combined self-supervised and vision–language objectives, nor has any research in the 3D imaging domain introduced a text generation objective.
3 Development framework
Due to the historical lack of public 3D vision–language datasets such as CT-RATE [12], the domain remains under-researched. We revisit key design decisions made by prior work to establish best practices for the 3D medical vision–language domain. This is conducted on chest CT as the region and modality of interest, due to the availability of a large image-only dataset (NLST) and a large paired image–report dataset (CT-RATE).
3.1 Pre-training datasets
CT-RATE
The CT-RATE dataset [12] consists of 25692 non-contrast CT acquisitions with associated reports from the Istanbul Medipol University Mega Hospital. Importantly, each report contains a Findings section, which describes the contents of the scan, and an Impression section, which represents an interpretation of the findings given the patient’s clinical history. Each acquisition in CT-RATE is expanded to 50188 unique 3D images by leveraging different reconstruction kernels. These kernels yield volumes with different spacings, with some reconstructions featuring high anisotropy, i.e., high in-plane resolution but low through-plane resolution, and others being more isotropic. As the reconstructions stem from the same image acquisition, their information content is highly redundant. Therefore, for each acquisition, we choose the reconstruction with the lowest in-plane size to minimise computational cost. This results in a median spacing of $\times0.7$$\times$1 mm and image size $\times512$$\times$359 voxels (distribution of in-plane sizes: 22417, 1648 and 43 images with size 512, 768, and 1024, respectively) for our subset of CT-RATE.
NLST
The NLST dataset [39] contains low-dose chest CT images from 26k patients, acquired at 33 different US centres, with each patient receiving one baseline scan and up to two follow-up scans with a one-year time difference between the scans, yielding a maximum of three scans per patient. Overall, this dataset provides about 72k unique 3D chest CTs without associated reports, with two reconstruction kernels each. Due to the two reconstructions being similar in spatial dimensions, we randomly pick a reconstruction kernel for each unique acquisition, yielding our subset of NLST.
As the images from both datasets have high overall dimensionality (median of $\times512$$\times$359 voxels for CT-RATE), we resample the images to 2 mm isotropic spacing to allow training our CLIP model with a FOV of the entire chest. For details on data preprocessing, we defer to Appendix A.
3.2 Global downstream tasks and datasets
To measure the quality and guide development of the trained vision and text encoder, we leverage global tasks, specifically image-to-report retrieval, classification probes, and zero-shot classification. We use CT-RATE to evaluate all of these tasks as it includes image-report pairs we use for image-to-report retrieval as well as multi-abnormality labels for the 18 most common abnormalities in the dataset, enabling zero-shot classification and the training of probes. To prevent data leakage between pre-training and downstream datasets, we divide the official CT-RATE training split into a train and a validation set, using the original splits employed in CT-CLIP (1k subjects with their associated reports and images) [12]. The official validation split of CT-RATE serves as a final test split.
During the development phase, we quantify retrieval performance through Recall at 1, 5, and 10 (R@1, R@5, R@10), and classification performance through AUPRC and AUROC, guiding the optimisation process. For a detailed explanation of the metrics, we refer to Sec. C.1.
For linear probing, we train five different sequence aggregation mechanisms with four different learning rates and a batch size of 16 for 15k steps with a cosine annealing learning rate schedule on the CT-RATE training dataset. The best performing probe is selected based on its performance on our split-off validation set from the training set of CT-RATE. This probe is later transferred for testing as-is to the test sets to yield the final prediction. More details on this are provided in Sec. C.2.
4 Methods and experiments
Translating the well-established CLIP method from the 2D imaging domain to the 3D medical domain is difficult due to the large domain gap and has been explored less due to the previously mentioned lack of publicly available data. In this section, we start from the basic CLIP paradigm, ablating various design choices, and iteratively extend the method with additional supervision objectives, yielding our comprehensive language–image pre-training (COLIPRI) encoder family (Fig. 2).
4.1 Vision–language 3D contrastive learning
Due to the large domain differences between natural imaging and 3D medical imaging, crucial training settings can vary substantially, requiring rediscovering well-tuned hyperparameters. To narrow the overall optimisation search space, we fix a few hyperparameters. Namely, we choose to train a Primus-M vision transformer (ViT) encoder [45] with an AdamW [22] optimiser. Each model is trained for 250k or 125k steps with a total batch size of 8 or 16, respectively, resulting in 2 million training samples being seen. We used 6.25k steps of linear learning rate warm-up, followed by a PolyLR schedule. We used a learning rate of $3\times 10^{-4}$ for batch size 8 and scaled it identically with the batch size . Aside from these fixed parameters, we chose an initial hyperparameter configuration that we optimise through a star sweep [1]. We choose a pre-trained BiomedVLP-CXR-BERT model [4] as the default text encoder due to the overlap of abnormalities between chest X-rays and chest CTs. By default, we pool dense vision and text tokens through a dedicated attention-pooling layer with 12 heads, trained from scratch for each encoder. We use the Findings section for supervision, and an input crop size of $\times192$$\times$192 at 2-mm isotropic spacing. Pre-training experiments are conducted on a single node 4 A100 (80 GB VRAM) GPUs unless specified otherwise.
Figure 3: Reports of 3D images are substantially longer than those of 2D images, leading to a distribution shift between long reports seen during training and short prompts used for zero-shot classification. Additionally, long reports might allow the text encoder to overfit due to high dimensionality instead of learning semantics.
4.1.1 Report length
Radiological reports often contain details about all organs imaged, stating whether the findings are normal or not. Should abnormalities be present, they are explicitly named; however, when absent, the abnormalities are often not listed, as this would yield an excessive list of abnormalities a patient does not suffer from. The exceptions are typically abnormalities that might have prompted the imaging study in the first place. This style of reporting results in reports of substantial sequence token lengths, with the average Findings section of CT-RATE being 243 tokens long when using the tokeniser of CXR-BERT (Fig. 3). This is in stark contrast to Zhang et al. [54] or Radford et al. [31], which either sample a single sentence from the paired text or whose datasets contain single sentence captions. Consequently, training with these long-form reports would inadvertently lead to a distribution shift when testing zero-shot classification with short-form prompts such as {abnormality} present”. Moreover, medical zero-shot classification is typically performed through negated statements (“No {abnormality} present.”), which may be a problem due to such statements being very sparse during training.
To account for this, we introduce two ways of conducting zero-shot classification: i) Native (N) zero-shot classification, where we average the embeddings of 50 long Findings sections from cases which are positive/negative for a particular abnormality. ii) Short (S) zero-shot classification, where we use the embeddings of {abnormality} present” and “No {abnormality} present” instead.
The overall shift between the two zero-shot prompting schemes is presented in Tab. 1, showing a 10 AUROC and AUPRC gap between them for our default CLIP configuration. The difference in performance between evaluation styles reveals that zero-shot classification in medical VLEs may be highly sensitive to linguistic formulation, with short diagnostic phrases (often not seen during training) yielding weaker alignment than native report-style embeddings.
Table 1: Report augmentation is important. Long reports allow overfitting of the text encoder. By introducing sentence shuffling and shortening through an LLM, this can be mitigated. Shuffle: Sentence Shuffle transform; Shorten: LLM sentence shortening.
Retrieval Probing Zero-shot (N) Zero-shot (S)
R@1 R@5 R@10 AUPRC AUROC AUPRC AUROC AUPRC AUROC
[…]
Shuffle 11.11 28.57 37.93 56.66 83.94 44.05 76.55 35.13 69.21
Shorten
p=[10%] 12.71 29.01 39.30 56.91 83.76 45.81 77.93 39.55 71.15
p=[25%] 11.54 27.68 38.04 56.45 83.87 47.12 78.67 35.13 68.10
p=[50%] 11.20 28.01 38.21 56.32 83.91 46.07 78.63 37.09 70.24
p=[75%] 9.78 25.92 34.36 56.97 84.01 46.41 78.74 34.70 68.19
To minimise the shift and reduce overfitting to the structure of long text reports, we introduce a Sentence Shuffle transform, which randomly shuffles the sentences of reports, substantially improving both retrieval and classification performance. In addition to sentence shuffling, we introduce a Short Sentence augmentation that replaces the long-form reports with a shortened version. These shortened reports were created using GPT-4333We used a gpt-4o (version 2024-08-06) Azure OpenAI endpoint to preprocess the reports.with instructions to reduce verbosity to a minimum. For details on the exact process, we defer to Appendix B. Combining Sentence Shuffle with the Short Sentence augmentations yielded further improvements in retrieval as well as classification performance (Tab. 1). In particular, the addition of the Short Sentence transform increases our simple zero-shot classification performance considerably, reducing the gap between our native and simple zero-shot classification settings. Aside from these two augmentations, less impactful aspects of the text were ablated, which are presented in Sec. B.2.
4.1.2 Field of view and number of patches
Table 2: Smaller patch-embedding is important, while having a global field of view is less relevant. To yield global representations, max-pooling performs very well for retrieval, while multi-head attention pooling performs well across the board. APE: Absolute Positional Encoding; Token Agg.: Token Aggregation.
Eval Retrieval Probing Zero-shot (N) Zero-shot (S)
Metric R@1 R@5 R@10 AUPRC AUROC AUPRC AUROC AUPRC AUROC
Patch Size
16x16x16 5.76 14.62 22.14 49.88 81.05 40.71 75.12 29.58 61.80
[…]
Input Size
128 8.35 22.06 29.66 56.19 83.78 44.60 77.17 32.00 63.05
160 7.94 23.64 32.75 55.91 83.51 43.22 76.24 34.77 66.70
[…]
224 - - - 54.44 83.06 42.69 76.13 29.99 64.78
no APE 9.27 23.48 33.33 55.38 83.20 43.30 76.18 25.07 58.56
Token Agg.
Avg Pool 5.93 18.88 27.90 54.34 83.03 40.20 75.48 35.65 67.46
Max Pool 11.45 27.57 38.01 56.18 83.56 42.80 75.32 25.86 56.57
SH-AP 5.43 19.13 27.57 54.05 82.82 40.84 75.83 33.27 62.19
[…]
Radiological reports typically describe findings across the entire image volume, rather than a restricted subregion. For example, chest CT reports primarily focus on pulmonary disease, but they also contain information about visible abdominal organs and other incidental findings. This means that many diagnostic tasks require a global FOV. In chest CT classification, the relevant abnormalities may be localised in specific lobes or distributed across the lungs. Therefore, the model must have access to the entire lungs FOV to avoid missing critical context. Training only on sub-crops may act as a form of regularisation, but, at inference time, such models may not classify images reliably without a global FOV since important abnormalities might lie outside the cropped region.
Practically, this requires very large input volumes for training. At a resolution of 2 mm isotropic spacing, an input size of $192^{3}$ voxels corresponds to a cube with edge 38.4 cm, which is sufficient to cover the lungs, which are commonly below 30 cm in all linear dimensions [18]. However, when using a patch size of $\times8$$\times$8 voxels in the ViT (the default in Primus, as larger patches often degrade performance on high-resolution dense downstream tasks [45]), the number of patches in a sequence reaches 14k tokens. This is orders of magnitude longer than typical vision–language settings, where natural images tokenised at standard patch sizes yield only 256 tokens [43].
This raises two fundamental questions: 1) Are large FOVs required for good performance? 2) How does one best aggregate this long token sequence into a global representation? To address these issues, we evaluate the effects of varying input size, token patch size, and token aggregation strategy, as well as removing absolute positional embeddings to allow varying the input size at test time, see Tab. 2.
Our results show that smaller input sizes are beneficial, while larger inputs reduce overall performance (Tab. 2 - Input Size). The only exception to this is the ‘short’ zero-shot classification, where an input size of $128^{3}$ performs worst. Smaller input sizes may improve performance by forcing the vision encoder to learn more robust and semantically meaningful representations. Larger FOVs expose all abnormalities simultaneously, allowing the encoder to rely on only a subset of correlated features. In contrast, smaller crops limit the visible context, incentivising the model to capture multiple discriminative cues. While this suggests that small input sizes as small as 128 may be better, an excessively small input size can’t capture the entire patient, limiting the applicability of the model. Additionally, we observe that reducing the sequence length through the use of a larger patch size has detrimental effects on overall performance (Tab. 2 - Patch Size), forcing us to keep the fine-grained tokens and rather long sequences at the cost of higher computational resources. Aggregating this sequence through max-pooling proved to be the best mechanism for retrieval and linear probing. However, our default multi-head attention pooling proved superior for zero-shot classification and yields competitive results across all metrics (Tab. 2 - Token Agg.), while allowing one to use the MaskCLIP trick [55] to generate language-aligned dense embeddings for segmentation. Lastly, we find that removing the absolute positional encoding (APE) only negatively affects short zero-shot segmentation, which is the least reliable metric. Hence, we chose to accept these minor penalties as a trade-off to allow dynamically adapting the input size (Tab. 2 - no APE).
4.1.3 Miscellaneous
Table 3: Miscellaneous hyperparameters. An optimal trade-off is achieved by balancing batch size with reduced training iterations. The sigmoid loss formulation does not improve performance. Lastly, minor spatial and not or low levels of intensity augmentation are best. Combining all changes of Sec. 4.1 yields our optimised CLIP model (COLIPRI-C), denoted in red, showing superior performance across all metrics.
Eval Retrieval Probing Zero-shot (N) Zero-shot (S)
Metric R@1 R@5 R@10 AUPRC AUROC AUPRC AUROC AUPRC AUROC
Batch Size - Training Steps
[…]
16 - 125k 9.77 24.98 32.75 55.66 83.44 41.68 75.48 31.82 63.40
24 - 62.5k 8.10 23.89 31.75 55.95 83.51 43.13 76.05 30.87 62.27
32 - 31.7k 8.27 20.05 29.32 55.19 83.23 40.55 74.77 23.69 53.89
[…]
Sigmoid Loss 5.60 16.21 24.06 53.88 82.81 39.45 74.66 30.90 63.44
Spatial - Intensity image augmentation
[…]
low - low 9.02 22.47 30.49 55.73 83.65 43.43 76.05 31.90 63.28
high - high 4.59 15.96 23.73 54.91 83.11 41.44 76.16 30.71 59.63
[…]
Contrastive learning in natural imaging benefits from large batch sizes, with e.g. 32k in  Tschannen et al. [43]. This is far out of reach for 3D medical imaging, where batch sizes are e.g., two when training segmentation models with nnU-Net [16], due to high VRAM consumption. Subsequently, we ablate the trade-off between larger batch sizes versus fewer iterations, while keeping the amount of seen samples identical. Lastly, we evaluate whether a sigmoid loss improves performance compared to the default softmax loss, and determine the extent to which strong spatial and intensity augmentations are necessary. Results are provided in Tab. 3. We find mid-sized batches with fewer iterations to be superior, observe a decrease in performance when using the sigmoid loss, and find low levels of intensity and spatial augmentations optimal, with stronger augmentations degrading performance.
4.1.4 Merging all changes
Based on our previous ablations, we introduce changes to our default configuration. We i) increase the batch size to 16, in conjunction with doubling our learning rate to $6\times 10^{-4}$ and halving our iterations to 125k, ii) we reduce the input size to $\times160$$\times$160, iii) add the Sentence Shuffle and Short Sentence text augmentation using LLMs, iv) add more image intensity augmentations, v) remove the absolute positional embedding as it is not necessary, and allows varying the input size of the model at test time. Results of the final, optimised CLIP model (COLIPRI-C) are presented in Tab. 3, showing improved performance across all metrics over the default configuration.
4.2 Including text generation
The goal of CLIP is to align image–report pairs. This objective can be a limiting factor in the medical domain, since there may exist multiple features that differentiate two image–report pairs, but a single one can suffice to distinguish them. This key insight spurred recent works to introduce the objective of predicting the image caption from the embedding of the image encoder [42, 47, 43]. To solve this task, the vision embeddings need to contain information about everything mentioned in the text report, as opposed to only about what differentiates two image–report pairs. Additionally, this objective is independent of the batch size, which is particularly important for a batch-size-constrained domain like 3D medical imaging. In this work, we combine the CLIP objective with a RRG objective based on CapPa [42], which conducts either causal captioning or parallel captioning in an interleaved fashion, i.e., alternating at each training iteration, to generate a report during training. Causal captioning refers to predicting the report in a next-word-prediction fashion using causal masking, while parallel captioning predicts the entire report simultaneously from a fully masked input.
4.2.1 Report generation for vision pre-training
As previously mentioned in Sec. 4.1, medical reports and natural image captions differ significantly, with medical reports being substantially longer and their structure more akin to a list. These aspects can pose hurdles in report generation, as the ordering of a listing is unpredictable without learning the preferences of the clinician who wrote the report or without memorising the entire report, both of which are undesirable. To address this, we use an LLM to structure the reports by assigning each sentence to one of eight semantic categories. Given these structured reports, we train our text decoder to generate the reports in a causal and a parallel fashion, but leave the section headers unmasked to guide the generation. This suffices for the captioning; however, for parallel captioning, we expect the amount of masked tokens between two section headers to leak information as no causal attention mask is used. This is due to sections being longer when pathological findings are present, which would allow the generative decoder to infer if diseases are present or not. To remove this bias, we group the headers at the start of the report, followed by a fixed amount of mask tokens. This informs the decoder of the desired ordering of the sections, without leaking the length of each section. For both of our generative tasks, we shuffle the order of the sections during training to regularise the decoder. Given this task formulation, we followed [42] and used a cross-attention-based approach to integrate vision tokens with a small transformer decoder. The generative and parallel text decoding setting is visualised in Fig. 4 with further details provided in Sec. C.4.
Figure 4: Text generation pre-training. To yield more semantic image features, we feed them through cross-attention into our text generation Eva02 transformer architecture, tasked with Captioning or Parallel Captioning. This is optimised simultaneously with the CLIP objective.
4.2.2 Report generation optimisations
Table 4: RRG objective optimisations. Shorter generator depth, lower RRG loss weight, and only using parallel captioning improve performance. Relative to the COLIPRI-C model, the inclusion of the RRG objective increases retrieval performance while reducing the classification performance. The optimised CLIP + RRG model (COLIPRI-CR), denoted in blue, shows superior performance across most metrics.
Retrieval Probing Zero-shot (N) Zero-shot (S)
R@1 R@5 R@10 AUPRC AUROC AUPRC AUROC AUPRC AUROC
Generator depth
4 9.52 25.31 33.25 55.85 83.40 46.29 76.53 37.69 69.41
6 11.70 26.57 35.42 55.72 83.28 44.73 75.88 34.56 69.39
8 10.44 24.56 32.41 54.14 82.59 45.20 75.80 31.23 60.90
[…]
CapPa-Cap Probability [%]
0-100 8.10 22.72 30.33 54.53 82.80 44.66 74.97 35.11 63.22
25-75 9.44 21.81 29.24 54.27 73.52 43.48 75.59 29.28 61.28
[…]
75-25 10.36 25.31 34.34 55.32 83.28 45.80 77.03 35.82 68.63
100-0 11.03 24.65 34.25 56.36 83.42 44.89 75.64 31.13 61.41
$\lambda_{\text{RRG}}$
0.1 11.03 26.15 35.09 55.79 83.14 45.56 76.43 33.65 64.59
0.3 9.27 24.48 33.00 55.29 83.17 45.38 76.70 36.46 68.09
[…]
When combining the optimised CLIP configuration (COLIPRI-C) with the RRG objective, there are various design decisions that warrant optimisation. By default, we choose a generator depth of 12 layers, a 50%-50% probability of causal versus parallel captioning, as well as a loss weight of $\lambda_{\text{RRG}}=1$, with $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CLIP}}+\lambda_{\text{RRG}}\cdot\mathcal{L}_{\text{RRG}}$. For these parameters, we conduct another star sweep from our default configuration, see Tab. 4. The results indicate that always using the parallel captioning loss is better than including only the causal captioning objective. This is likely due to the next-word captioning objective learning to memorise long reports, resulting in the vision tokens not being required for the decoding. Moreover, we observed that lower decoder depths are more beneficial than deeper decoders, forcing vision embeddings to be quickly adoptable for report generation. Lastly, we observe an intermediate loss weight of 0.3 to be optimal. Combining these changes yields our optimised CLIP + CapPa model (COLIPRI-CR), see at the bottom of Tab. 4. Interestingly, we observe the additional generative decoding objective to substantially increase Retrieval performance; however, it does not translate to increases in linear probing or zero-shot classification, even incurring slight decreases in linear probe and native zero-shot classification and moderate reductions in simple zero-shot classification performance. This indicates that the current generative decoding objective may still be subject to confounders that prevent the vision encoder from learning semantically meaningful representations, which would allow linear separation between pathological abnormalities.
4.3 Including vision-only self-supervision
Table 5: MAE optimisations. Increasing the masked autoencoder decoder depth to 6, using block masking, and including the MAE loss only for the last 25% of training improves performance. Other changes were deemed not relevant. Optimised CLIP + MAE model (COLIPRI-CM), denoted in green.
Retrieval Probing Zero-shot (N) Zero-shot (S)
R@1 R@5 R@10 AUPRC AUROC AUPRC AUROC AUPRC AUROC
MAE Decoder
2 14.04 30.91 39.77 55.15 82.99 44.61 76.31 38.15 70.84
[…]
6 14.45 33.42 41.52 55.71 83.44 45.22 77.04 38.14 71.50
8 12.78 29.99 39.68 55.04 82.95 44.73 76.81 37.30 70.70
Masking Ratio
60% 13.78 30.83 38.85 55.08 83.17 44.71 76.14 38.76 72.57
[…]
90% 14.20 32.08 41.85 55.41 83.02 44.32 76.32 37.92 71.94
Mask Style
[…]
Block 14.87 29.74 38.68 55.46 83.12 44.63 76.84 39.33 72.45
Inverse Block 13.37 30.41 40.02 55.44 83.19 45.01 76.83 39.22 72.37
Included at last [X%] of training
25% 13.95 29.07 36.51 56.41 83.83 45.42 76.57 41.99 74.95
50% 12.95 29.91 38.68 56.40 83.39 45.26 76.63 40.65 72.83
75% 14.54 30.08 38.35 55.62 83.28 45.64 77.47 36.92 70.81
[…]
$\lambda_{\text{MAE}}$
0.1 14.62 29.16 37.59 54.99 83.07 44.82 76.57 32.75 66.17
0.5 13.62 29.91 38.68 55.34 83.06 44.77 76.49 40.10 72.25
[…]
2.0 14.62 31.75 40.52 55.08 83.07 45.05 76.96 37.31 69.74
Minimal isotropic spacing included
[…]
1 mm 15.5 31.7 39.8 56.19 83.44 43.16 76.48 41.35 74.90
0.5 mm 14.3 29.6 38.6 55.19 82.98 43.38 76.12 40.67 73.51
[…]
While image–text pre-trained encoders tend to learn useful representations for global reasoning tasks, they require paired data and their learned representations are often less powerful for dense tasks [27], which represent the majority of the challenges in the medical imaging community [45]. To improve the quality of the learned embeddings of our vision encoder for dense tasks, we pair our language–image pre-training with an additional MIM, vision-only objective. Given the recent OpenMind benchmark [44], we choose MAE as our vision-only objective, mostly due to the lack of more advanced dense pre-training methods such as iBOT, DINOv2, and DINOv3 in 3D medical imaging. To simplify integration of the vision-only objective, we integrate the MAE pre-training with our COLIPRI-C, postponing the final CLIP, RRG, and MAE integration to the end.
MAE objectives are prevalent in the domain, and often subject to high masking ratios of 60-90% [25, 46]. These high masking ratios occlude the majority of the image, which we expect to complicate the CLIP objective. To minimise this interference, we alternate between vision-only and vision–language objectives for each training batch, following [43]. Moreover, due to the vision-only objective not being subject to the FOV problem (Sec. 4.1), we can expose our models to sub-crops at higher resolutions and incorporate NLST into our training data.
4.3.1 Vision-only optimisations
Analogous to RRG, the inclusion of the MAE introduces various factors of variation deserving ablation. In particular, we ablate the masking ratio, masking style (random, block and inverse block masks), mask decoder depth, the vision-only loss weight $\lambda_{\text{vo}}$, as well as ablating if later inclusion of the vision-only objective is beneficial. Moreover, we assess the inclusion of higher-resolution images, evaluating the effect of including images resampled to 1-mm and 0.5-mm isotropic spacings. Results are visualised in Tab. 5.
We observe it to be beneficial to include the vision-only objective later in the training of the vision–language model. We find the masking ratio of 75% to be optimal in combination with random masking, as well as using an equal vision-only loss weight of 1. Additionally, we encountered training stability issues when training with the same learning rate, so it was reduced by a factor of 2 to $3\times 10^{-5}$.
4.3.2 CLIP, RRG and MAE
Table 6: COLIPRI development results. Compared to the COLIPRI-C objective COLIPRI-CRM shows increases in retrieval, but decreases in linear probing and native zero-shot classification performance. This is not fully surprising as the added objectives do not focus solely on aligning the global embeddings.
Retrieval Probing Zero-shot (N) Zero-shot (S)
R@1 R@5 R@10 AUPRC AUROC AUPRC AUROC AUPRC AUROC
[…]
Having integrated the MAE with CLIP and RRG with CLIP, we merge the optimal configurations (see Secs. 4.2.2 and 4.3.1) of CLIP + RRG and CLIP + MAE without further ablations, yielding our final COLIPRI-CRM method. The final validation results are displayed at the bottom of Tab. 5. The inclusion of both MAE and RRG objectives yields a slight decrease in linear classification probe performance and native zero-shot classification performance relative to COLIPRI-C configuration. This is partially to be expected due to the MAE objective function focusing on the quality of dense embeddings, which were not quantified. A more holistic evaluation is provided in Sec. 6.
5 Experiments
We evaluate our optimised COLIPRI encoders in Tab. 6 on multiple unimodal (semantic segmentation, multilabel classification) and multimodal (zero-shot classification, report generation) tasks.
5.1 Classification and report generation
Classification performance is evaluated on the withheld test set of CT-RATE and additionally on the publicly available subset of RAD-ChestCT [10], which comprises 3.6k chest CT volumes with 16 multi-abnormality labels that can be derived from the original CT-RATE abnormality classes. Similarly, as during development, we evaluate linear classification probes, as well as native and short zero-shot classification performance on both aforementioned datasets.
As an additional global task, we evaluate the quality of the frozen image encoder embeddings for report generation. To do this, we follow the LLaVA framework [21], with image tokens passed through a two-layer multilayer perceptron (MLP) to integrate them into the language space of the Qwen 2.5 1B base model [41]. We focused on generating the Findings section of each report. To evaluate the clinical accuracy of generated reports, we use the text classifier trained by Hamamci et al. [12] based on RadBERT [51] as well as RadFact-CT (+/-) and (+), variants of RadFact [2] with CT-specific system prompts and few-shot examples, see Sec. C.1 for details. Subsequently, both clinical and lexical metrics are calculated on the CT-RATE test set.
As baselines, we compared our method against established CT models, namely CT-CLIP [12], and CT-FM [28], all trained on chest CT datasets; Merlin [3], which, despite being trained on abdominal CTs, we find to be a competitive baseline for chest CT reporting.
5.2 Semantic segmentation
To evaluate the quality of the vision encoder for dense tasks, we measure 3D medical image segmentation performance after fine-tuning the encoder. In this setting, we compare against a Primus-M[45] encoder trained from scratch, as well as a MAE-pre-trained [44] Primus-M encoder trained on our CT-RATE and NLST data (1mm and 2mm isotropic spacings). All training runs are conducted using the nnU-Net framework [16]. All of the encoders are fine-tuned for 37.5k steps using nnU-Net [16], following the best, short, Primus-M training rate schedule determined in Wald et al. [44], but with peak learning rate reduced to $1\times 10^{-4}$. To remain partially in distribution, we chose to focus on segmentation datasets with targets in the upper abdomen or chest region, a FOV that is often visible during pre-training. On each segmentation dataset, we train the first three folds of a five-fold cross-validation. As datasets, we choose
  1. 1.
    LiTS [36] ($N$ = 131), task 3 of the Medical Segmentation Decathlon (MSD), contains segmentations for liver and liver tumours, often still within the FOV of chest CTs.
  2. 2.
    Lung [36] ($N$ = 64), task 6 of the MSD, which contains cases of primary lung cancers.
  3. 3.
    KiTS23 [14] ($N$ = 489), a dataset focused on segmenting tumors, cysts and the kidney.
On all datasets, we report the Dice similarity coefficient (DSC), averaged across all foreground classes.
6 Results and discussion
6.1 Classification probes
Table 7: Classification probing results. We compare the embedding quality of our pre-trained vision encoders against publicly available baselines. In particular, our COLIPRI-C model yields the best classification results, exceeding all baselines on both datasets across all metrics. Notably, our classification pipeline yields notably higher classification values. The metrics in curly brackets are from in [33]. Differences in performance with the metrics we computed using the released checkpoints may be due to configuration issues. Hence, we report both sets of values for fairness and clarity. AUPRC: area under precision- recall curve; AUROC: area under receiver operating characteristic curve; BA: balanced accuracy; $F_{1}$: (non-weighted) $F_{1}$-Score.
CT-RATE RAD-ChestCT
AUPRC AUROC BA $F_{1}$ AUPRC AUROC BA $F_{1}$
CT-CLIP 25.96 61.21 {75.1} 57.66 {67.6} 34.49 28.77 54.05 {64.7} 52.65 {62.5} 39.08*
CT-FM 53.54 82.14 73.51 55.56 42.41 68.49 61.95 47.55
Merlin 54.81 82.62 74.28 56.69 45.30 70.91 64.34 49.35
[…]
We evaluate the quality of our VLEs and other baseline vision encoders through classification probes on CT-RATE and RAD-ChestCT (Tab. 7).
Compared to reference values from literature, we observe that our probing setup is largely superior to previous probe setups, yielding AUROC values of above 80% for the majority of baseline methods (vs. approx. 75% reported in the original works) as well as our own encoders. This likely originates from our training setup leveraging multiple token aggregation schemes as well as multiple learning rates for each of the encoders. Moreover, not all of our token aggregation schemes are linear as some include a light-weight non-linear attention pooling block, which is more flexible than a linear layer. However, we believe this multi-probe scheme to be more suited for comparison due to its robustness to hyperparameter selection, and generally increases the performance for all encoders. The only exception for this performance increase is the CT-CLIP encoder, which curiously performs worse in our experiments. This might be due to a potential configuration issue, hence we additionally report the results from Shui et al. [33] for the in-common metrics.
Our COLIPRI models exceed all baseline methods across all evaluated metrics, with COLIPRI-C representing the strongest pre-training method for classification. In particular, for the non-thresholded metrics AUPRC and AUROC, it increases by 2 points and 1.5 points, respectively, over the best-performing baseline Merlin. [^4] Merlin being the best baseline is commendable as it was originally trained on abdomen CTs instead of chest CTs. However, it proved itself a strong baseline not only for classification probing but also for zero-shot classification, as reported in Shui et al. [33].
The inclusion of our RRG and MAE objectives decreases classification probe performance slightly, which is likely due to the added objectives not being focused on improving the linear separability of abnormalities.
6.2 Zero-shot classification
Table 8: Zero-shot classification results. Comparing the zero-shot capability, we observe that our CLIP encoder using ‘native’ prompts performs competitively to fVLM, without requiring segmentation masks at inference. Additionally, while our encoders exceed the remaining baselines with the ‘native’ prompting scheme, the performance degrades a lot when using ‘short’ prompts. An extended version of this table with additional metrics is provided in Tab. 13. PS: prompt style; BA: balanced accuracy; $F_{1}: w$eighted-$F_{1}$ score.
CT-RATE RAD-ChestCT
Model PS AUROC BA w-$F_{1}$ AUROC BA w-$F_{1}$
CT-CLIP - 70.4 65.1 69.1 63.2 59.9 64.8
BIUD - 71.3 68.1 71.6 62.9 60.6 65.2
Merlin - 72.8 67.2 70.9 64.4 61.9 66.3
fVLM - 77.8 71.8 75.1 68.0 64.7 68.8
[…]
*Values from [33], which excluded ‘Medical Material’ and ‘Lymphadenopathy’.
For all aligned vision–language encoders, we report the zero-shot classification performance on CT-RATE and RAD-ChestCT, using the ‘short’ and ‘native’ prompting schemes (Sec. 4.1.1) for our models (Tab. 8). For our baselines, we leverage some of the values from the fVLM paper [33]. However, in their work, results are reported for only 16 of the 18 abnormalities in CT-RATE, excluding the non-localisable ‘Lymphnodeadenopathy’, and ‘Medical Material’ as they are not associated with an organ [^5] https://github.com/alibaba-damo-academy/fvlm/issues/12#issuecomment-3283463870, which is a limitation of fVLM.
The ‘native’ prompting scheme is still substantially better than the ‘short’ prompting scheme. Our COLIPRI-C configuration decreases the most with 7 points, while COLIPRI-CM decreases the least with about 3 points in AUROC on CT-RATE. This indicates that our text encoder does not embed the short-form prompts in a semantically similar fashion as when averaging the embeddings of the long-form reports. Unfortunately, we could not evaluate the effect of the different prompting styles on the baselines, but we expect them to reveal that CT-CLIP and BIUD would benefit much more from a ‘native’ prompting scheme, as they were similarly trained to our encoders – by aligning embeddings of long-form reports. On the other hand, we expect Merlin and fVLM to benefit less, if at all, as they decompose their reports into shorter report sections for different regions-of-interest, which exposes them less to long-form reports and decreases the distribution shift to short prompts.
Our COLIPRI-C encoder using ‘native’ prompting performs similarly to fVLM on CT-RATE, reaching an AUROC of 77.8% and 75.2% w-$F_{1}$. Transferred to RAD-ChestCT, it decreases 11 points to 66.98% AUROC, and performs slightly worse than fVLM, which decreases 10 points. COLIPRI-CRM performs worse on the CT-RATE dataset, reaching 75.02% AUROC, but generalises to RAD-ChestCT, yielding similar performance with ‘native’ prompts as COLIPRI-C. This difference in performance between CT-RATE and RAD-ChestCT is shared for all our pre-trained encoders that include additional loss objectives aside from the CLIP objective. We hypothesise that this is due to the regularising effects of the additional loss objectives, which reduce overfitting to dataset-specific features and encourage the encoder to learn more domain-invariant representations that generalise better across CT datasets.
6.3 Report generation
[…]
Lexical report metrics
ROUGE-L 54.2 [53.0, 55.3] 52.4 [51.4, 53.4] 53.2 [52.1, 54.3] 54.8 [53.6, 55.9] 54.8 [53.6, 56.0] 54.6 [53.4, 55.8] 54.8 [53.6, 56.0]
BLEU-1 55.6 [54.3, 57.1] 57.8 [56.6, 59.1] 58.7 [57.5, 59.9] 61.8 [60.8, 63.0] 62.5 [61.4, 63.6] 61.9 [60.9, 63.1] 61.6 [60.5, 62.8]
BLEU-4 41.0 [39.7, 42.4] 41.2 [40.0, 42.6] 42.2 [40.9, 43.5] 44.2 [42.9, 45.5] 44.5 [43.1, 45.9] 44.3 [43.0, 45.7] 44.1 [42.8, 45.4]
METEOR 53.7 [52.5, 54.8] 53.5 [52.5, 54.6] 54.6 [53.6, 55.7] 56.8 [55.7, 57.8] 57.1 [56.0, 58.3] 56.7 [55.6, 57.8] 56.8 [55.7, 57.9]
Clinical report metrics
[…]
RadFact-CT/Logical $F_{1}$ (+/-) 68.7 [67.7, 69.6] 67.4 [66.4, 68.4] 68.2 [67.3, 69.2] 70.3 [69.3, 71.3] 70.5 [69.4, 71.5] 70.0 [68.9, 71.0] 69.2 [68.1, 70.2]
RadFact-CT/Logical $F_{1}$ (+) 12.8 [11.4, 14.2] 11.5 [10.5, 12.7] 11.7 [10.8, 12.7] 18.0 [16.7, 19.2] 20.4 [19.2, 21.7] 20.0 [18.8, 21.4] 19.4 [18.4, 20.8]
RadFact-CT/Logical Precision (+) 14.8 [13.0, 16.8] 13.9 [12.6, 15.4] 14.1 [12.9, 15.4] 21.5 [20.1, 23.2] 24.9 [23.2, 26.7] 24.3 [22.8, 25.9] 23.6 [22.0, 25.2]
RadFact-CT/Logical Recall (+) 11.3 [9.7, 12.7] 9.9 [8.9, 11.0] 10.1 [9.1, 11.1] 15.4 [14.1, 16.8] 17.3 [16.1, 18.6] 17.0 [15.8, 18.4] 16.5 [15.5, 17.9]
We evaluate the impact of our pre-training strategy on downstream radiology report generation using both lexical and clinical metrics (Fig. 5, Tab. 9). Across all lexical metrics – ROUGE-L, BLEU1, BLEU4, and METEOR – our pre-trained encoders perform on par with or slightly above the baselines, indicating that all methods allow generating similarly lexically accurate reports. However, our clinical metrics show clearer distinctions between methods, confirming that high lexical overlap does not necessarily imply clinically accurate reports. This is consistent with recent work in chest x-ray report generation [19], which highlights that lexical metrics do not reliably correlate with clinical correctness.
[…]
6.4 Segmentation
We evaluate semantic segmentation performance on the LiTS, Lung and KiTS23 datasets, against a baseline trained from scratch as well as an MAE pre-trained on NLST and CT-RATE using the same images at 1-mm and 2-mm isotropic resolutions, using nnssl [44] (Tab. 10).
MAE pre-training exceeds both the baseline trained from scratch and our pre-trained COLIPRI encoders, confirming its place as the best pre-training method for dense downstream tasks. Despite our encoders focusing on a contrastive pre-training objective, which was shown to struggle with exceeding a baseline trained from scratch for segmentation tasks in Wald et al. [44], our encoders yield slight improvements in DSC on the LiTS and KiTS datasets over training from scratch.
Surprisingly, the inclusion of the MAE objective into the pre-training regime has minor effects on the segmentation results, generally reaching a lower average performance than the model trained with the RRG objective. We hypothesise this to be a result of the combination of the late-stage inclusion of the MAE objective, as proposed in [43] and determined to be optimal for global classification performance in Sec. 4.3.
Table 10: Segmentation fine-tuning results. DSC results of a Primus-M 3D ViT trained for 150 epochs (37.5k steps) from scratch, from a pre-trained MAE initialisation, and from our configurations. Across all segmentation tasks, the MAE pre-trained encoder performs best, while our encoders yield better results than from scratch on LiTS and KiTS23, despite including contrastive pre-training objectives. Fold $k$: fold $k$ of a cross-validation set of runs.
Pre-training LiTS Lung KiTS23
Fold 0 Fold 1 Fold 2 mean Fold 0 Fold 1 Fold 2 mean Fold 0 Fold 1 Fold 2 mean
From scratch 78.24 71.03 74.27 74.51 72.10 53.60 61.78 62.49 79.34 78.96 77.49 78.60
MAE 81.79 78.96 81.50 80.75 74.03 57.26 65.03 65.44 85.03 84.09 84.24 84.45
[…]
Runs with unexpectedly low performance.
(a) Input
(b) Merlin
(c) CT-FM
(d) CT-CLIP
(e) COLIPRI-C
(f) COLIPRI-CM
(g) COLIPRI-CR
(h) COLIPRI-CRM
[…]
6.5 Qualitative analysis
Aside from quantitative results, we provide a principal component analysis (PCA) of the embeddings of Merlin, CT-FM, CT-CLIP, and our COLIPRI encoders, on a lung cancer case from the MSD Lung dataset (Fig. 6).
The resolution of the embeddings of Merlin and CT-FM is very low, providing hardly any localisation of semantics. CT-CLIP yields embeddings of higher resolution, allowing one to map the features from the input chest CT to the PCA map. However, the PCA is inconsistent and noisy, and shows high sensitivity to air in its principal components, and a strong bias towards position embeddings. On the other hand, our COLIPRI encoders yield higher-resolution embeddings, which are sharper and more consistent, allowing for clear recognition of the boundaries of the patient, lungs, and the abdominal organs, as well as the lung mass present in the right lung (on the left-hand side of the coronal and axial slice views).
6.6 Summary of the results
Our encoders exhibit strong linear separability of features to abnormality classes, outperforming previous baselines and confirming that contrastive pre-training yields discriminative and semantically rich embeddings, as demonstrated by our classification probing results (Sec. 6.1). In the zero-shot classification setting (Sec. 6.2), however, performance depends heavily on the prompt formulation: models perform well with ‘native’ prompts resembling the training distribution but underperform when ‘short’ prompts are used, suggesting a high sensitivity to prompting style.
In report generation (Sec. 6.3), our models enable the creation of more comprehensive and factually consistent reports, as reflected by higher RadFact and RadBERT scores relative to the baselines. This indicates that our vision encoder representations effectively capture clinically relevant information from imaging data.
Segmentation results (Sec. 6.4), in turn, lag far behind MAE performance and show minimal improvement for some configurations. This suggests that combining the global alignment objective with the masked image encoder objective does not synergise well to enhance fine-grained spatial feature quality without hindering the learning of global semantics. Overall, our findings demonstrate that our approach effectively strengthens global vision–language alignment for holistic medical understanding, while leaving performance on dense tasks largely unaffected.
Across all our experiments, we find that our COLIPRI encoders excel at global tasks, particularly in report generation and classification probing, while showing no change or slight improvements over models trained from scratch for semantic segmentation.
7 Limitations and conclusion
Overall, our encoders substantially strengthen the global alignment between vision and language features, producing strong performance in classification investigations and report generation. However, there remain several important limitations and opportunities for improvement.
Firstly, the sensitivity to prompt formulation in a zero-shot setting underscores limited robustness and generalisability under prompt shifts (‘native’ vs ‘short’). Mitigating this brittleness may require strategies such as prompt augmentation, prompt contrastive training, or architectures that support prompt adaptation or prompt-invariant representations.
Secondly, the limited impact of our pre-training on segmentation performance suggests that combining global alignment and MIM with a low-level voxel reconstruction objective may not be complementary. Voxel-wise reconstruction primarily emphasises local appearance fidelity rather than high-level semantics; this can lead to a mismatch with the contrastive alignment objective, which operates in the embedding space. In contrast, more advanced MIM strategies, such as iBOT’s online tokeniser and embedding-level consistency objective [56], might offer a stronger synergy between the pre-training paradigms. By encouraging the student network to match teacher embeddings for masked tokens, such approaches preserve both spatial structure and global semantic coherence, potentially yielding richer representations relevant for dense downstream tasks such as segmentation.
Nonetheless, our results show that contrastive pre-training in a vision–language paradigm is a promising direction for holistic 3D medical image understanding. These methods generate richer semantic alignment between images and text, which directly benefits tasks such as classification, retrieval, and report generation.
Appendix A Data preprocessing
A.1 Image preprocessing
To conduct pre-training and classification experiments, we preprocessed CT-RATE, NLST and RAD-ChestCT to a unified format using TorchIO [29]. This preprocessing consisted of
  1. 1.
    Reorientation of all images to RAS+ image orientation.
  2. 2.
    Dividing the CTs Hounsfield units by 1000, effectively mapping -1000 to -1 and +1000 to 1, followed by clipping values outside of this range.
  3. 3.
    Resampling to 2-mm, 1-mm and 0.5-mm isotropic spacing (1 mm and 0.5 mm were only used in conjunction with the vision-only pre-training paradigm) using an antialiasing filter for downsampling, and B-Spline interpolation.
CT-RATE contains various Head CT images which were removed before pre-training. CT-RATE and NLST contain images derived from the same acquisition, yielding redundant information. For CT-RATE, we only keep the reconstruction with the lowest spacing; for NLST, we keep one randomly selected image from each acquisition, as their spacings are largely similar.
A.2 Report preprocessing
CT-RATE contains image–report pairs, with each report containing Findings and Impression sections, as well as other sections we did not use. The reports in the released dataset were originally translated from Turkish to English using the Google Translate API We processed the reports using GPT-4o as explained below.
We re-translated the reports (Sec. E.1), structured the Findings (Sec. E.2) and split the sections (Sec. E.3) into short sentences of positive and negative findings for a certain anatomical region or semantic topic. The sections are 1) Image Quality 2) Lungs and Airways 3) Pleura 4) Mediastinum and Hila 5) Cardiovascular Structures 6) Bones and Soft Tissues 7) Tubes, Lines, and Devices 8) Upper Abdomen. Given these sections, reports are processed in two ways:
  1. 1.
    Each sentence gets assigned in its original, long state to one of these sections, effectively structuring the Findings (prompt provided in Sec. E.2). Below is an example of the Lungs and Airways section:
    The trachea and both main bronchi are patent, with no obstructive pathology detected. Ventilation of both lungs is normal, and no mass or infiltrative lesion is observed. Additionally, there is a hypodense lesion measuring 15 mm in diameter located in the posterolateral middle segment of the left lung, possibly a cyst.
  2. 2.
    Similar to 1., each sentence is assigned to a region, shortened, and classified as a positive finding (i.e., mentions the presence of an abnormality) a positive finding (i.e., mentions the absence of a abnormalities). An example of the latter structure is displayed in Fig. 8 for the Bones and Soft Tissues section, and the prompt used to create these is provided in Sec. E.3.
Appendix B Optimising CLIP hyperparameters
B.1 Language augmentations
Due to the issues arising from the long-form reports of 3D radiological images, we introduce two text data augmentations aimed at reducing overfitting and improving our CLIP models’ short-form zero-shot classification performance using short-form statements.
Figure 7: Sentence shuffling regularises the order of sentences, removing potential ordering biases of practitioners when writing their reports.
B.1.1 Sentence shuffling
The length of reports in CT-RATE allows the text encoder to overfit easily, e.g., by learning to distinguish subjects through their unique sequence of tokens. In order to force our text encoder to learn more semantic patterns, we introduce an augmentation that breaks apart the sentences of our report by splitting it at the ”.” delimiter and randomly permuting the sentences. While this may seem rigorous, the majority of medical reports reflect a listing, which does not really follow a consistent order. Even more importantly, this augmentation removes biases of e.g., practising radiologists, which have an implicit bias in how they prefer inspecting and reporting on an image. By introducing this augmentation, this batch effect can be removed, forcing our model to focus on the semantics instead. An example of report shuffling is given in Fig. 7.
B.1.2 Sentence shortening
Figure 8: Sentence shortening reduces the domain shift between long reports seen during training time and short texts for simple zero-shot classification.
Table 11: Evaluating the influence of various changes to the text used for training our CLIP model. Impr.: Training with impressions instead of findings; Find. + Impr.: Training with findings and impressions appended to the findings; Re-translate: Using a recent GPT endpoint to re-translate the reports from Turkish to English; Shuffle: Using the Sentence Shuffle augmentation; DnC: Did not Converge;
Eval Retrieval Probing Zero-shot (N) Zero-shot (S)
Metric R1 R5 R10 AUPRC AUROC AUPRC AUROC AUPRC AUROC
[…]
Sentence Shuffle 11.11 28.57 37.93 56.66 83.94 44.05 76.55 35.13 69.21
[…]
Impressions 7.77 21.80 30.58 54.89 83.29 43.74 76.52 29.11 62.40
Find. + Impr. 8.69 22.22 31.83 55.31 83.27 43.64 76.74 31.39 67.64
[…]
Re-translate 7.53 20.90 29.35 55.24 83.00 42.69 76.29 28.29 60.11
Re-translate + Shuffle 9.95 24.58 32.27 56.21 83.64 44.22 76.73 32.05 67.19
BiomedCLIP 3.43 13.28 19.47 52.79 81.89 37.43 74.21 28.16 59.30
CXR-BERT (scratch) DnC DnC DnC DnC DnC DnC DnC DnC DnC
[…]
During training, the text encoder is only exposed to reports of substantial length. However, when conducting zero-shot classification, a user may prefer to query with very brief, single sentence statements. This shift between training and test time can lead to substantial performance differences because the text encoder has never been trained on such data. To reduce this shift, we create abbreviated, structured reports using GPT-4o (Fig. 8). We instruct GPT to minimise verbosity and distinguish between pathological (positive) findings and statements about healthy anatomy (negative) findings. Given these shortened sentences, we concatenate them in arbitrary order to formulate our new, shortened findings, replacing the original, long reports with a certain probability. We ablate various probabilities of applying this transformation and ablate two versions of this augmentation, one where we only create findings with positive statements and one where we use both positive and negative findings. Results of our positive + negative findings are visualised in Tab. 12. We observe that lower probabilities increase the zero-shot classification performance, while increasing amounts of this augmentation leads to slight increases in probe performance. While this augmentation was aimed at improving simple zero-shot classification performance, we found zero-shot classification to be one of the noisiest metrics. Due to the AUROC of the probing increasing with higher probability and the increases in ‘native’ zero-shot AUROC, we chose p=[25%] as the application probability of this transformation in our COLIPRI-C model.
Table 12: Results of the Sentence Shortening augmentation for different probabilities. All results are in conjunction with sentence-shuffling to remove the confounding effect of the random concatenation order.
Eval Retrieval Probing Zero-shot (N) Zero-shot (S)
Metric R1 R5 R10 AUPRC AUROC AUPRC AUROC AUPRC AUROC
[…]
p=[10%] 12.71 29.01 39.30 56.91 83.76 45.81 77.93 39.55 71.15
p=[25%] 11.54 27.68 38.04 56.45 83.87 47.12 78.67 35.13 68.10
p=[50%] 11.20 28.01 38.21 56.32 83.91 46.07 78.63 37.09 70.24
p=[75%] 9.78 25.92 34.36 56.97 84.01 46.41 78.74 34.70 68.19
B.2 Additional text ablations
Aside from investigating text augmentations, we evaluate the influence of training with impressions or findings and impressions, as well as training with a ’better’ translation of the original Turkish reports to English, aimed at improving clinical lingo. Results are visualised in Tab. 11, highlighting that findings and impressions are better than only impressions; however, using only the Findings is superior to both. Moreover, results show that the newer translation does not positively affect performance.
B.2.1 Text encoder
Our default text encoder is BiomedVLP-CXR-BERT, a transformer pre-trained on reports of chest X-rays. Since this model is rather small, holding about 110M parameters, and was trained on substantially shorter X-ray reports, we chose to evaluate the effect of using a larger text encoder, namely the 196M parameter large BiomedCLIP model pre-trained on the entire PubMed collection, as well as training our CXR-BERT architecture from scratch. Results are presented in Tab. 11 at the bottom, showing the superiority of BiomedVLP-CXR-BERT over the larger BiomedCLIP model.
Appendix C Evaluation details
C.1 Metrics
C.1.1 Retrieval metrics
Recall
Given image report-pairs, we evaluate the image-to-report retrieval through the Recall @ 1/5/10. This is calculated by embedding the entire validation set or test set of image-report pairs, yielding e.g. 1292 validation images and reports (The value is larger than 1000 due to some patients having multiple sessions – head CT images are removed though).
This yields 1292 global image embeddings and 1292 report embeddings, for which we calculate the similarities between all pairs, identically as during the CLIP training. Following this, we measure whether the image embedding of the actual image is the most similar, within the five most or within the 10 most similar images. From these results, we compute Recall@1, Recall@5, and Recall@10, which quantify the proportion of test samples for which the correct image appears among the top-1, top-5, or top-10 retrieved results, respectively. A higher recall value indicates that the learned embedding space more effectively aligns visual and textual representations, allowing relevant image–report pairs to be retrieved more reliably.
C.1.2 Classification metrics
Area under the receiver operating characteristic curve (AUROC)
The AUROC metric evaluates a model’s ability to distinguish between positive and negative cases across all possible decision thresholds and is computed as the area under the curve defined by the True Positive Rate (sensitivity) plotted against the False Positive Rate (1–specificity). An AUROC of 0.5 indicates random performance, whereas a value of 1.0 represents perfect discrimination.
Area under the precision-recall curve (AUPRC)
The AUPRC metric measures a model’s ability to identify positive cases across varying decision thresholds, emphasising performance on imbalanced datasets. It is computed as the area under the curve defined by Precision (positive predictive value) plotted against Recall (sensitivity). Unlike AUROC, which considers both positive and negative classes equally, AUPRC focuses on the model’s effectiveness in detecting the positive class, making it particularly informative when positive cases are rare, as is the case for many abnormalities. A higher AUPRC indicates that the model maintains strong precision even at high recall levels, reflecting its capacity to identify true positives while minimising false detections.
Due to this, we determine the best probe as the probe that yields maximal AUPRC on the validation set.
$F_{1}$-score ($F_{1}$)
To compute the $F_{1}$-score, a decision threshold must be defined to distinguish predicted positives from predicted negatives. For each abnormality, this threshold is selected as the value that maximises the $F_{1}$-score on the internal CT-Rate validation split. Once determined, the threshold remains fixed for all subsequent evaluations on the test sets. In contrast to the original CT-CLIP study [12], we report a non-weighted $F_{1}$-score, as the unweighted metric reflects the model’s ability to classify individual abnormalities more accurately. We use the term $F_{1}$ score synonymously for the macro $F_{1}$ score, unless explicitly specified. We only resort to using the weighted $F_{1}$ score when comparing baseline values we were not able to run.
Balanced accuracy (BA)
Balanced Accuracy measures a model’s overall classification performance while accounting for class imbalance, yielding 0.5 for random chance and 1.0 for perfect accuracy. We calculate the balanced accuracy by reusing the same decision boundary optimised for the $F_{1}$-score, even though it may not be the optimal threshold to maximise BA.
C.1.3 Report generation metrics
RadBERT
RadBERT [12] is a text classification BERT model trained on CT-RATE, which allows to predict class probabilities for the 18 different multi-abnormality classes of the CT-RATE dataset. Through it we evaluate the report generation quality of the encoders quantitatively through Micro and Macro $F_{1}$-scores.
RadFact (+/-)and RadFact (+)
RadFact, originally proposed by Bannur et al. [2], is a metric that assesses the factuality of each sentence in a generated report, by evaluating if the sentence is supported by a reference (ground-truth) report. This is achieved by leveraging the reasoning capabilities of GPT-4o.
Because our data differ from the X-ray reports used in the original work, we adapt RadFact’s system prompt and introduce two distinct RadFact variants: RadFact (+/-) and RadFact (+). RadFact-CT (+/-) evaluated both positive and negative radiological statements, while RadFact-CT (+) focuses exclusively on positive findings – excluding statements about the absences of abnormality, unremarkable observations, or normal anatomy.
In this study, we employ RadFact’s Logical Precision and Logical Recall to compute a Logical $F_{1}$ score. The grounding and spatial reasoning capabilities of RadFact are not considered in our evaluation.
C.2 Classification linear probing
Given a pre-trained encoder we conduct linear-probing to measure the quality of our vision encoders embedding for classifying the abnormalities labelled in CT-RATE. To do so, we discard the original token aggregation scheme of the vision encoder, which was aimed at aligning image and report, and instead train a new one for classification. Due to not knowing which Token aggregation scheme is best, we conduct a grid-search over five different schemes and four different learning rates. The token aggregation schemes are as follows:
  1. 1.
    Average Pooling: A simple averaging across the sequence dimension.
  2. 2.
    Max Pooling: A simple max-pooling across the sequence dimension.
  3. 3.
    Learned Attention Pooling: An attention pooling head with a learned query token, steering how the tokens are recombined to yield the final global representation.
  4. 4.
    Average Attention Pooling: Same as above, just with the learned query replaced by a token created through average pooling.
  5. 5.
    MultiLearnedAttentionPool: The same as Learned Attention Pooling, just with four query tokens instead of one. As we get one representation for each query, the four outputs are averaged to yield the global representation.
All of these token pooling schemes yield a global embedding of embedding dimension size, which we consequently project down through a linear layer to the 18 abnormalities annotated in the CT-Rate dataset. The four learning rates we sweep are $lr\in\{10^{-1},3\cdot 10^{-2},1\cdot 10^{-2},3\cdot 10^{-3}\}$. Due to keeping the encoder frozen, we can allocate the majority of VRAM to the probes, allowing us to train all of them jointly, as in [27]. The training itself is conducted using a batch size of 16, for 12.5k steps using an SGD optimiser with momentum 0.95 and 0 weight decay following a cosine annealing learning rate schedule. Due to the input volumes being larger than the input size, we conduct center-cropping of the volume, extracting an e.g. central $160\times 160\times 160$ crop (the same as the vision encoders input size). Once training concluded, the single best probe is selected based on the area under the precision-recall curve (AUPRC) values on the validation set. For thresholded metrics, we select a unique threshold for each multi-abnormality class based on the optimal $F_{1}$ score. This is necessary as we train our probe with a binary cross-entropy loss, which doesn’t take the imbalance of the multi-abnormalities into account, yielding a decision boundary that is offset from 0.5. When using the probes for testing, the same probe with their associated thresholds are translated to the test set yielding the final metrics.
C.3 Zero-shot classification
Opposed to the trained classification probes of Sec. C.2 the originally trained multi-head attention pooling as well as the language-encoder and language-pooling is reused to evaluate zero-shot classification performance. In this paper we differentiate between two zero-shot classification schemes.
  1. 1.
    Native (N): For native zero-shot classification, we aggregate 50 reports of a patient with an abnormality and 50 reports of patients without this abnormality. Each of the long Findings is passed through the language encoder and the language pooling to yield 50 embeddings for positives and 50 embeddings for negatives. Each group is averaged, to yield a representative embedding of the abnormality being present or absent from the reports.
  2. 2.
    Short (S): For each abnormality, a small template is used to create sentences about whether an abnormality is present or absent. In our case, we use {abnormality} present’ ‘no {abnormality} present’. The resulting embedding represents the presence or absence of this abnormality.
Given the language embeddings representing the presence or absence of an abnormality, a global vision embedding is extracted from a centre-crop of each image. For each of these global vision embedding the cosine similarity between the vision and the two language embeddings is calculated, the similarities are temperature-scaled (divided by 0.07), and the resulting logits are fed through a softmax to yield probabilities associated with the presence and absence of the abnormality. The probability associated with the positive embedding is used to calculate the threshold-less and thresholded metrics as in Sec. C.2.
C.4 Report generation
We tested the potential of our patch embeddings on the vision–language task of report generation. For the language component, we used the Qwen2.5-1B base model [40], which was not instruction-tuned to ensure fair evaluation of intrinsic alignment.
Our training recipe adheres to the LLaVA-style framework [21], where a canonical frozen vision encoder and trainable decoder paradigm is used for multimodal vision–language generation. The 3D vision backbone remains frozen throughout training to preserve pre-learned visual representations. On top of this encoder, we train both a cross-modal alignment module and the language decoder. We employ a causal language modelling loss with teacher forcing, applied to tokenised radiology reports. The optimisation objective is thus purely autoregressive, and no auxiliary objectives are introduced.
Following prior work [21, 30], we integrated vision tokens into the language space through a two-layer MLP projection head. We constrain supervision to the Findings section of the CT report. The Findings provide high-density, structured clinical interpretation of the CT volume, covering organ-level abnormalities and radiographic evidence. In contrast, the Impression section, although often used in clinical practice, introduces redundancy without providing additional information that could be extracted from the input image. We therefore omit it in all experiments.
Each vision–language model (VLM) is fine-tuned on the CT-RATE training set with a batch size of 32 for 10 epochs with no weight decay. The maximum learning rate is $5\times 10^{-5}$ and a cosine learning rate schedule is used with a linear warm-up for 3% of the training steps. These hyperparameters were selected to maximise the Micro-$F_{1}$ scores on the validation set.
C.5 Image-to-report retrieval
Image-to-report retrieval evaluates how well a model aligns visual and text representations in a shared embedding space. By retrieving the correct clinical report given an image, we directly measure whether the model captures clinically meaningful visual semantics and associates them with corresponding textual descriptions. This task thus serves as a strong proxy for multimodal understanding and vision–language alignment. Due to the objective of pre-training being image-report alignment, no additional adaptation step is required for this task. Hence, the vision and text encoders with their respective pooling mechanisms are used as-is to evaluate this task.
C.6 Segmentation fine-tuning
To evaluate segmentation performance, we leverage the pre-training adaptation framework proposed in nnssl [44], which introduces fine-tuning of pre-trained vision encoders into the well-established nnU-Net framework[16]. In particular, a longer training schedule of 1000 nnU-Net epochs (250k iterations) and a shorter training schedule of 150 nnU-Net epochs (37.5k iterations) were proposed in this paper. We leverage both to evaluate the embedding quality of our vision encoders, with details on the explicit settings available in Wald et al. [44] and the nnssl repository [^6] https://github.com/MIC-DKFZ/nnssl.
Dataset preprocessing
During pre-training we trained our vision encoder on CT data that was rescaled from -1000/+1000 to -1/+1 and clipped to -1/+1. However, in initial tests, we found this to yield sub-par results when using it for semantic segmentation. Consequently, we stick to the official nnU-Net normalisation, referred to as ‘CTNormalisation’, which clips values outside the 0.5th percentile and the 99.5th percentile before standardising to zero-mean and unit variance (standardisation is conducted on the dataset and not the image level). Moreover, despite the majority of encoders trained on 2-mm isotropic spacing (with the exception of some MAEs), we chose to resample the segmentation datasets to 1-mm isotropic spacing, as this resolution is substantially closer to the median spacings the downstream datasets come with. We note that this shift in normalisation and spacing is not optimal and negatively influences segmentation performance. While the spacing issue is not easily avoidable, the normalisation choice could be adapted easily in future work.
Appendix D Additional results
We provide additional zero-shot classification results in Tab. 13.
Table 13: Additional zero-shot classification results complementing Tab. 8. PS: prompt style; AUPRC: area under precision recall curve; AUROC: area under receiver operating characteristic curve; BA: balanced accuracy; $F_{1}: w$eighted $F_{1}$ score; Values taken from Shui et al. [33], where the abnormalities ‘Lymphadenopathy’ and ‘Medical material’ were excluded.
Model PS CT-RATE RAD-ChestCT
AUPRC AUROC BA $F_{1}$ w-$F_{1}$ AUPRC AUROC BA $F_{1}$ w-$F_{1}$
CT-CLIP - - 70.4 65.1 - 69.1 - 63.2 59.9 - 64.8
BIUD - - 71.3 68.1 - 71.6 - 62.9 60.6 - 65.2
Merlin - - 72.8 67.2 - 70.9 - 62.9 60.6 - 65.2
fVLM - - 77.8 71.8 - 75.1 - 68.0 64.7 68.8 -
[…]
Appendix E Prompts
E.1 Prompt to translate reports from Turkish to English
"""You are a board-certified radiologist-translator.
Translate the Turkish radiology report contained inside a single <report>  </report> element into fluent, precise English.
##  OUTPUT  COPY THIS SHAPE EXACTLY
**1. Clinical Information**
English text here.
**2. Technique**
English text here.
**3. Findings**
English text here.
**4. Impression**
English text here.
 **The four numbered headings must stay exactly as above and remain in bold.**
 If any section is empty, whitespace, or literally nan, write Not provided. (plain text, **not** bold) under that heading.
 Do **NOT** output anything outside these four labelled sections.
 No bullet characters (, , *, etc.) or markdown lists inside the body text.
##  INPUT
You will receive one well-formed XML block:
<report>
  <clinical_information></clinical_information>
  <technique></technique>
  <findings></findings>
  <impression></impression>
</report>
##  STYLE RULES
 Literal, complete translation  no omissions, additions, or summaries.
 Concise, objective radiology tone (passive voice preferred).
 Use RSNA / ACR terminology; convert decimal commas to periods (7,5 mm  7.5 mm).
 Expand abbreviations on first mention: CT pulmonary angiography (CTPA).
 Preserve original sentence order and punctuation.
##  REQUIRED GLOSSARY  replace the Turkish term with the English term verbatim
buzlu cam grnts              ground-glass opacity
plevral efzyon                  pleural effusion
septal kalnlama                interlobular septal thickening
konsolidasyon                    consolidation
akcier nodl                   pulmonary nodule
retikler opasiteler             reticular opacities
broniektazi                     bronchiectasis
hiler lenfadenopati              hilar lymphadenopathy
mediastinal ift                 mediastinal shift
trakea orta hatta                trachea is midline
perikardiyal efzyon             pericardial effusion
pheli kitle                    suspicious mass
subplevral bant                  subpleural band
havayolu duvar kalnlamas      airway wall thickening
lenf bezi bymesi               lymph-node enlargement
ateli infiltrasyon              inflammatory infiltration
atelektazi                       atelectasis
bal petei grnm              honeycombing pattern
fibrotik deiiklikler           fibrotic changes
amfizem                          emphysema
tomurcuklanm aa              tree-in-bud pattern
kontrastsiz                      non-contrast enhanced
kontrast verilmeden              non-contrast enhanced
E.2 Prompt to structure Findings sections into different subsections
"""You are a radiology report editor.
Restructure a non-contrast chest-CT report (supplied in four free-text blocks)
into the fixed template below **without altering a single medical fact**.
  INPUT
The incoming text always uses these bold labels:
**Clinical Information:**
**Technique:**
**Findings:**
**Impression:**
  OUTPUT
Copy this skeleton exactly. Section and subsection titles must be **bold** and end
with a colon. After each colon insert one space, then the content or the fallback line.
**1. Clinical Information:**
**2. Technique:**
**3. Comparison:**
   If prior imaging referenced; else: No prior imaging available for comparison.
**4. Findings:**
**4.1 Image Quality:**
   If no limitations: Diagnostic image quality. No significant artifacts noted.
**4.2 Lungs and Airways:**
   If no pulmonary findings: No pulmonary abnormalities detected.
**4.3 Pleura:**
   If no pleural findings: Pleura unremarkable.
**4.4 Mediastinum and Hila:**
   If no findings: Mediastinal and hilar structures unremarkable.
**4.5 Cardiovascular Structures:**
   If no findings: Cardiovascular structures unremarkable.
**4.6 Bones and Soft Tissues:**
   If no findings: No osseous or soft-tissue abnormalities detected.
**4.7 Tubes, Lines, and Devices:**
   If none present: No tubes or devices identified.
**4.8 Upper Abdomen:**
   If unremarkable or not imaged: No upper-abdominal abnormalities detected.
**5. Impression:**
   If missing: No impression provided.
  EDITING RULES
 Zero-omission: every medical statement from the original Findings and Impression
  MUST reappear once (and only once) in an appropriate subsection.
 Do not add, delete, combine, or reinterpret abnormalities.
 Re-phrase into concise, passive radiology English (RSNA/ACR style).
 If a section/subsection is entirely absent, insert the exact fallback line.
 No lists, bullets, metadata, or commentaryreturn only the final formatted report.
After drafting, mentally cross-check that every clinical phrase from the original is present.
Begin when you receive the four-block input.
E.3 Prompt to extract positive and negative findings
You are an AI assistant that makes radiology reports more succinct. These reports are being used
to train a 3D CLIP-style deep learning model. You will be given the full findings section.
You will extract, for each of the 8 sections in the findings text, a list with negative findings
and a list with positive findings.
In the first list, you must collect a summarized sentence for each negative finding mentioned.
For example, a sentence like "Esophagus is within normal limits.  In the sections passing through
the upper part of the abdomen, the bilateral adrenal glands appear natural. No significant
pathology was detected in the abdominal sections." must be mapped to a list like  ["Normal
esophagus.", "Natural bilateral adrenal glands.", "No abdominal pathologies."]. The exact sentences
must be short but maintain their core message. Positive findings are not allowed in this list
and have to be ignored.
In the second list you must summarize only the positive findings that are denoted.
In this version sentences like ’The heart and mediastinal vascular structures have
a natural appearance’, ’Esophagus is within normal limits.’, ’No occlusive pathology
was detected in the trachea and both main bronchi.’ or ’Trachea and main bronchi are open.’
have to be left out. When positive (pathological) findings are mentioned, summarize them
very briefly. E.g. a sentence like ’atypical infiltration areas of septal thickenings are
observed in places’ can be summarized as ’Septal thickenings.’. Similarly as before create
a list of short sentences about positive abnormalities ["Septal thickenings.", "Multiple lung
nodules.", ...]. Make sure the sentences you create are a statement and less of a description,
like how someone would search for the case as opposed to how one would describe it
in a findings report.
Ignore all information that cannot possibly be predicted from the corresponding single image
or provided clinical information section. Any comparison or reference to prior imaging must
be ignored from the output. Do not output findings about how the image was acquired.
Output this in JSON format with one key for each of the eight sections. Each section is a mapping
from section name (e.g. "image quality" or "cardiovascular structures") to the "negative findings"
and "positive findings" lists. This is the structure:
  "image_quality": {
    "negative_findings": [
    "positive_findings": [
  "lungs_and_airways": {